OpenStudy (anonymous):

An alloy of tin is 15% tin and weighs 20 pounds. How many pounds of tin are in the alloy?

OpenStudy (anonymous):

0.15*20 = 3

OpenStudy (anonymous):

An alloy of tin is 15% tin and weighs 20 pounds. A second alloy is 10% tin. How many pounds of the second alloy must be added to the first to get a 12% mixture?

OpenStudy (anonymous):

Solving (0.15*20+0.10*x)/(x+20) = 0.12*(x+20) 3+0.1x=0.12(x^2+40x+400) 0.12x^2+4.7x+45=0

OpenStudy (anonymous):

I might be wrong though... I'm not so sure now.